# BAD-NEUS: Rapidly converging trajectory stratification

### v2

@@ Abstract
An issue for molecular dynamics simulations is that events of interest often involve timescales that are much longer than the simulation time step, which is set by the fastest timescales of the model. Because of this timescale separation, direct simulation of many events is prohibitively computationally costly. This issue can be overcome by aggregating information from many relatively short simulations that sample segments of trajectories involving events of interest. This is the strategy of Markov state models (MSMs) and related approaches, but such methods suffer from approximation error because the variables defining the states generally do not capture the dynamics fully. By contrast, once converged, the weighted ensemble (WE) method aggregates information from trajectory segments so as to yield unbiased estimates of both thermodynamic and kinetic statistics. Unfortunately, errors decay no faster than unbiased simulation in WE. Here we introduce a theoretical framework for describing WE that shows that introduction of an approximate stationary distribution on top of the stratification, as in nonequilibrium umbrella sampling (NEUS), accelerates convergence. Building on ideas from MSMs and related methods, we generalize the NEUS approach in such a way that the approximation error can be reduced systematically. We show that the improved algorithm can decrease simulation times required to achieve a desired precision by orders of magnitude.
@@ I Introduction
Conceptually, one can view our approach as a form of stratification, in which the weights of not just individual trajectory segments but groups of them are manipulated. Stratification was first introduced for controlling the sampling of trajectory segments in nonequilibrium umbrella sampling (NEUS) [3, 16, 17, 4, 18, 19]. NEUS groups trajectories that are in the same regions of state space, as defined by collective variables (CVs), and estimates the steady-state probabilities of the regions by solving a global flux balance equation; this strategy was subsequently adopted in extensions of WE [20] and exact milestoning (EM)[6]. A unified framework for trajectory stratification that incorporates elements of all of the above algorithms and is unbiased in the limit that each region contains an infinite number of ensemble members is presented in Ref. 19, which also shows that the regions can be defined in terms of path-based quantities. Mathematical analysis of trajectory stratification algorithms can be found in Refs. 21, 22.
@@ II Weighted Ensemble (WE)
Here, we use a relatively simple procedure and represent the splitting and pruning by resampling the ensemble within each region. Each iteration of the sampling thus consists of two steps: evolution according to the underlying dynamics and resampling the ensemble. In the evolution step, for each walker $i$, we numerically integrate for a time interval $\Delta$ to obtain $X^{i}_{t+\Delta}$ from $X^{i}_{t}$ and update $J^{i}_{t+\Delta}$ accordingly. In this step the weights are unchanged; $w^{i}_{t+\Delta}=w^{i}_{t}$. We note that $\Delta$ can be a random stopping time, not just a fixed time interval. An often useful choice is to take $t+\Delta$ to be the first time after time $t$ that the value of the index process $J^{i}$ changes. In this case, each walker has a different value for $\Delta.$ In the resampling step, for each value $k$ of the index process (e.g., for each region in a partition of state space), if there is at least one walker with $J^{i}_{t+\Delta}=k$, we select a number $N_{k}$. Then we sample from the set $\Phi=\{i:J^{i}_{t+\Delta}=k\}$ $N_{k}$ times with replacement according to the probabilities $p_{i}=w^{i}_{t+\Delta}/\sum_{i\in\Phi}w^{i}_{t+\Delta}$. For each index $r$ so chosen, we append $(X^{r}_{t+\Delta},J^{r}_{t+\Delta},\sum_{i\in\Phi}w^{i}_{t+\Delta}/N_{k})$ to the new ensemble.
@@
$\mathbbm{E}_{X_{0}\sim\pi}[g(X_{0},X_{1},\ldots,X_{\Delta-1})]\\ =\lim_{T\to\infty}\frac{1}{T}\sum_{n=0}^{T}\frac{1}{N}\sum_{i=0}^{N}g(X^{i}_{n\Delta},X^{i}_{n\Delta+1},\ldots,\\ X^{i}_{n\Delta+\Delta-1})w^{i}_{n\Delta},$ (1)
@@
$\overline{\pi}_{\ell}(dx|k)=P[X_{S(\ell)}\in dx|J_{S(\ell)}=k]$ (2)
where $dx$ is an infinitesimal volume in state space that is located at a specific value of $x$, and $S(\ell)$ is an increasing sequence of stopping times. For example, if we choose $S(\ell)=\ell\Delta$ for a fixed $\Delta$, the following description corresponds to the standard version of WE already sketched. Alternatively, we can let $S(\ell)$ be the time of the $\ell$-th change in the value of the index process $J$:
@@
$\overline{z}_{\ell}^{k}$ $=P[J_{S(\ell)}=k]\approx\sum_{i=1}^{N}w_{S^{i}(\ell)}^{i}\mathbbm{1}_{\{k\}}\left(J_{S^{i}(\ell)}^{i}\right).$ (5)
@@
$\sum_{j}\int g(x,j)\left(\frac{1}{\zeta_{k}}\mathbbm{1}_{\{k\}}(j)\rho(dx,j)\right)=\\ \mathbbm{E}\left[\sum_{j}\int g(x,j)\mathcal{R}_{k}\rho\right]$ (12)
@@
$\mathbbm{E}_{(x,k)\sim\pi}[g(X_{0},J_{0},\dots,X_{t},J_{t})]=\frac{\sum_{k}\overline{z}^{k}\int\overline{\pi}(dx|k)\mathbbm{E}_{X_{0}=x,J_{0}=k}\left[\sum_{t=0}^{S(1)-1}g(X_{t},J_{t},\dots,X_{t+\tau},J_{t+\tau})\right]}{\sum_{k}\overline{z}^{k}\int\overline{\pi}(dx|k)\mathbbm{E}_{X_{0}=x,J_{0}=k}\left[S(1)\right]},$ (16)
@@
1: Approximate Markov propagator $\tilde{\mathcal{U}}$, resampling operator $\mathcal{R}_{k}$, starting flux distributions $\{\overline{\pi}_{0}(dx|k)\}_{k=1}^{n}$, initial region weights $\{\overline{z}_{0}^{k}\}_{k=1}^{n}$ with $\sum_{k}\overline{z}_{0}^{k}=1$.
@@
9:return $\overline{\pi}_{L}(dx|k)$ and $\overline{z}_{L}^{k}$
@@ III Nonequilibrium umbrellla sampling (NEUS)
1: Approximate Markov propagator $\tilde{\mathcal{U}}$, Resampling operator $\mathcal{R}_{k}$, starting flux distributions $\{\overline{\pi}_{0}(dx|k)\}_{k=1}^{n}$, initial region weights $\{\overline{z}_{0}^{k}\}_{k=1}^{n}$ with $\sum_{k}\overline{z}_{0}^{k}=1$.
@@
11:return $\overline{\pi}_{L}(dx|k)$ and $\overline{z}_{L}^{k}$
@@
If the flux distributions and region weights assume their steady-state values such that $\overline{\pi}_{\ell}(dx|k)=\overline{\pi}(dx|k)$ and $\overline{z}_{\ell}^{k}=\overline{z}^{k}$, a global balance condition is satisfied:
@@ IV BAD-NEUS
$0=\sum_{k}\overline{z}^{k}\int\overline{\pi}(dx|k)\mathbbm{E}_{X_{0}=x,J_{0}=k}\left[\sum_{t=0}^{S(1)-1}f(X_{t},J_{t})-f(X_{t+\tau},J_{t+\tau})\right].$ (26)
@@
$0=\sum_{k}\sum_{p}c_{\ell}^{p}\overline{z}_{\ell}^{k}\int\phi_{p}(x,k)\overline{\pi}_{\ell}(dx|k)\mathbbm{E}_{X_{0}=x,J_{0}=k}\left[\sum_{t=0}^{S(1)-1}\phi_{r}(X_{t},J_{t})-\phi_{r}(X_{t+\tau},J_{t+\tau})\right].$ (28)
@@
$M_{pr}=\sum_{k}\overline{z}_{\ell}^{k}\int\phi_{p}(x,k)\overline{\pi}_{\ell}(dx|k)\mathbbm{E}_{X_{0}=x,J_{0}=k}\left[\sum_{t=0}^{S(1)-1}\phi_{r}(X_{t},J_{t})-\phi_{r}(X_{t+\tau},J_{t+\tau})\right].$ (30)
@@
1: Approximate Markov propagator $\tilde{\mathcal{U}}$, Resampling operator $\mathcal{R}_{k}$, starting flux distributions $\{\overline{\pi}_{0}(dx|k)\}_{k=1}^{n}$, initial region weights $\{\overline{z}_{0}^{k}\}_{k=1}^{n}$ with $\sum_{k}\overline{z}_{0}^{k}=1$, basis set $\{\phi_{p}(x,k)\}_{p}$, lag time $\tau$, total number of iterations $L.$
@@
13:return $\overline{\pi}_{L}(dx|k)$ and $\overline{z}_{L}^{k}$
@@
1: Starting ensemble of walkers $E_{0}=\{(X_{0}^{i},J_{0}^{i},w_{0}^{i})\}_{i=1}^{N},$ with $\sum_{i=1}^{N}w_{0}^{i}=1$, number of past iterations to retain $h$, total iterations $L,$ number of strata $n.$ lag time $\tau$
@@
1: List of trajectories $T_{\ell}$, lag time $\tau$, basis functions $\phi_{p}$, bin weights $\overline{z}_{\ell}$
@@
When working with walkers, this mixture distribution corresponds to concatenating the walkers from $h$ prior iterations, and therefore contains more data than using a single iteration. We simply substitute $\overline{\pi}_{\ell}$ for $\overline{\pi}_{\ell,h}$ in all of our algorithms. A unified walker-based algorithm that we use in practice is given in Algorithm 4, with the additional steps needed for NEUS and BAD-NEUS in Algorithm 5.
@@ V.1 Comparison of algorithms
Figure 2: Convergence of the steady-state distribution of the Müller-Brown system for different algorithms as indicated.
@@ V.2 Choice of hyperparameters
Having demonstrated that trajectory stratification outperforms WE for this system, we now examine the effect of key hyperparameters in NEUS and BAD-NEUS.
@@ V.3 Kinetic statistics
Thus there are $n+m$ total strata, the first $n$ of which contain walkers that last visited $A$, and the remainder of which contain walkers that last visited $B$. The steady-state distribution of walkers which last visited $A$ is proportional to $\pi(x)q_{-}(x)$, and that which last visited $B$ is proportional to $\pi(1-q_{-}(x))$. This ensemble lets us compute dynamical averages. The backward committor projected onto a space of CVs can be computed using
@@ VI Molecular Test System
All molecular dynamics simulations were performed with OpenMM [41]. We compute CVs including backbone RMSDs and fractions of native contacts using MDtraj [42]. To enable direct comparison with Ref. 37, we model NTL9(1-39) using the AMBER FF14SB force field [43] and the Hawkins, Cramer and Truhlar generalized Born implicit solvent model [44, 45]. We use a Langevin thermostat with a time step of 2 fs, a temperature of 300. K, and a friction constant of $\gamma=80$ ps-1, which corresponds to the high-viscosity case considered in Ref. 37.
@@
where $Q(X)$ is the fraction of native contacts using the definition of Ref. 46. We determine the native contacts from the crystal structure, PDB ID 2HBB [47]. We space 20 $Q_{k}$ values uniformly in the interval $[0.35,0.85]$ and set $\varepsilon_{k}=0.6(Q_{k+1}-Q_{k})$, so that the regions overlap. Thus there are 40 total strata. We define the folded state as full backbone RMSD $<0.28$ nm and $Q>0.85$; the denatured state has full backbone RMSD $>0.80$ nm and $Q<0.35$.
@@
We run each walker until its index process changes and then an additional 10 ps, so the trajectory length and time per iteration are random variables. For the NTL9 simulations, the average trajectory length was 50 ps, and the average time per iteration was 4 $\mu$s. The results in Figure 9 thus required about 204 $\mu$s in total, comparable to the 252 $\mu$s for the haMSM-WE calculation. The time to compute the rate constant by BAD-NEUS and haMSM-WE is much less than the millisecond timescale of a single folding event, which we also expect to be the time to converge WE without any acceleration strategy. BAD-NEUS actually provides some speedup relative to haMSM-WE in the sense that it provides potentials of mean force and committors in its present form, whereas a calculation in the unfolding direction (performed either separately or simultaneously, as here) would be required to obtain these statistics from haMSM-WE. By the same token, we likely could reduce the computation time for BAD-NEUS by “teleporting” walkers from $B$ to $A$ and computing only the forward rate [50]. However, we expect most investigators would seek the mechanistic information provided by the potentials of mean force and committor (which can be used to also obtain reactive currents [24]) if they are investing the computational resources to compute the rate, and free energies computed from ratios of forward and backward rates can serve as checks on the rates [51, 52].
@@
Figure 10: (top) BAD NEUS bin weights for the NTL9(1-39) system. Strata with indices less than 20 correspond to the folding direction (trajectories last in the denatured state), and strat with indices greater than or equal to 20 correspond to the unfolding direction (trajectories last in the native state). (bottom) Deviation of weights ($\ln(\overline{z})$) from those in the last iteration.

### v3

@@ Abstract
An issue for molecular dynamics simulations is that events of interest often involve timescales that are much longer than the simulation time step, which is set by the fastest timescales of the model. Because of this timescale separation, direct simulation of many events is prohibitively computationally costly. This issue can be overcome by aggregating information from many relatively short simulations that sample segments of trajectories involving events of interest. This is the strategy of Markov state models (MSMs) and related approaches, but such methods suffer from approximation error because the variables defining the states generally do not capture the dynamics fully. By contrast, once converged, the weighted ensemble (WE) method aggregates information from trajectory segments so as to yield unbiased estimates of both thermodynamic and kinetic statistics. Unfortunately, errors decay no faster than unbiased simulation in WE as originally formulated and commonly deployed. Here we introduce a theoretical framework for describing WE that shows that introduction of an approximate stationary distribution on top of the stratification, as in nonequilibrium umbrella sampling (NEUS), accelerates convergence. Building on ideas from MSMs and related methods, we generalize the NEUS approach in such a way that the approximation error can be reduced systematically. We show that the improved algorithm can decrease simulation times required to achieve a desired precision by orders of magnitude.
@@ I Introduction
Conceptually, one can view our approach as a form of stratification, in which the weights of not just individual trajectory segments but groups of them are manipulated. Stratification was first introduced for controlling the sampling of trajectory segments in nonequilibrium umbrella sampling (NEUS) [3, 16, 17, 4, 18, 19]. NEUS groups trajectories that are in the same regions of state space, as defined by collective variables (CVs), and estimates the steady-state probabilities of the regions by solving a global flux balance equation; this strategy was subsequently adopted in extensions of WE [20] and exact milestoning (EM)[6]. A unified framework for trajectory stratification that incorporates elements of all of the above algorithms and is unbiased in the limit that each region contains an infinite number of ensemble members is presented in Ref. 19, which also shows that the regions can be defined in terms of path-based quantities. Mathematical analyses of trajectory stratification algorithms can be found in Refs. 21, 22.
@@ II Weighted Ensemble (WE)
Here, we use a relatively simple procedure and represent the splitting and pruning by resampling the ensemble within each region (or, more generally, value of the index process). Each iteration of the sampling thus consists of two steps: evolution according to the underlying dynamics and resampling the ensemble. In the evolution step, for each walker $i$, we numerically integrate for a time interval $\Delta$ to obtain $X^{i}_{t+\Delta}$ from $X^{i}_{t}$ and update $J^{i}_{t+\Delta}$ accordingly. In this step the weights are unchanged; $w^{i}_{t+\Delta}=w^{i}_{t}$. We note that $\Delta$ can be a random stopping time, not just a fixed time interval. An often useful choice is to take $t+\Delta$ to be the first time after time $t$ that the value of the index process $J^{i}$ changes [3, 16, 17, 4, 18, 19]. In this case, each walker has a different value for $\Delta.$ In the resampling step, for each value $k$ of the index process (e.g., for each region in a partition of state space), if there is at least one walker with $J^{i}_{t+\Delta}=k$, we select a number $N_{k}$. Then we sample from the set $\Phi=\{i:J^{i}_{t+\Delta}=k\}$ $N_{k}$ times with replacement according to the probabilities $p_{i}=w^{i}_{t+\Delta}/\sum_{i\in\Phi}w^{i}_{t+\Delta}$. For each index $r$ so chosen, we append $(X^{r}_{t+\Delta},J^{r}_{t+\Delta},\sum_{i\in\Phi}w^{i}_{t+\Delta}/N_{k})$ to the new ensemble.
@@
$\mathbbm{E}_{X_{0}\sim\pi}[g(X_{0},X_{1},\ldots,X_{\Delta-1})]=\lim_{T\to\infty}\frac{1}{T}\sum_{n=0}^{T}\frac{1}{N}\sum_{i=0}^{N}g(X^{i}_{n\Delta},X^{i}_{n\Delta+1},\ldots,X^{i}_{n\Delta+\Delta-1})w^{i}_{n\Delta},$ (1)
@@
$\overline{\pi}_{\ell}(dx|k)=\mathbbm{P}[X_{S(\ell)}\in dx|J_{S(\ell)}=k]$ (2)
where $dx$ is an infinitesimal volume in state space that is located at a specific value of $x$, and $S(\ell)$ is an increasing sequence of stopping times. For example, if we choose $S(\ell)=\ell\Delta$ for a fixed $\Delta$, the following description corresponds to the standard version of WE already sketched. Alternatively, we can let $S(\ell)$ be the time of the $\ell$-th change in the value of the index process $J$, as in NEUS[3, 16, 17, 4, 18, 19],
@@
$\overline{z}_{\ell}^{k}$ $=\mathbbm{P}[J_{S(\ell)}=k]\approx\sum_{i=1}^{N}w_{S^{i}(\ell)}^{i}\mathbbm{1}_{\{k\}}\left(J_{S^{i}(\ell)}^{i}\right).$ (5)
@@
$\int g(x,k)\left(\frac{1}{\zeta_{k}}\rho(dx,k)\right)=\mathbbm{E}\left[\int g(x,k)\mathcal{R}_{k}\rho\right]$ (12)
@@
$\mathbbm{E}_{(X_{0},J_{0})\sim\pi}[g(X_{0},J_{0},\dots,X_{\tau},J_{\tau})]=\frac{\sum_{k}\overline{z}^{k}\int\overline{\pi}(dx|k)\mathbbm{E}_{X_{0}=x,J_{0}=k}\left[\sum_{t=0}^{S(1)-1}g(X_{t},J_{t},\dots,X_{t+\tau},J_{t+\tau})\right]}{\sum_{k}\overline{z}^{k}\int\overline{\pi}(dx|k)\mathbbm{E}_{X_{0}=x,J_{0}=k}\left[S(1)\right]},$ (16)
@@
1: Approximate Markov propagator $\tilde{\mathcal{U}}$, resampling operator $\mathcal{R}_{k}$, starting flux distributions $\{\overline{\pi}_{0}(dx|k)\}_{k=1}^{n}$, initial region weights $\{\overline{z}_{0}^{k}\}_{k=1}^{n}$ with $\sum_{k}\overline{z}_{0}^{k}=1$, total number of iterations $L$, number of strata $n$.
@@
9:return $\{\overline{\pi}_{L}(dx|k)\}_{k=1}^{n}$ and $\{\overline{z}_{L}^{k}\}_{k=1}^{n}$
@@ III Nonequilibrium umbrellla sampling (NEUS)
1: Approximate Markov propagator $\tilde{\mathcal{U}}$, Resampling operator $\mathcal{R}_{k}$, starting flux distributions $\{\overline{\pi}_{0}(dx|k)\}_{k=1}^{n}$, initial region weights $\{\overline{z}_{0}^{k}\}_{k=1}^{n}$ with $\sum_{k}\overline{z}_{0}^{k}=1$, total number of iterations $L$, number of strata $n$.
@@
11:return $\{\overline{\pi}_{L}(dx|k)\}_{k=1}^{n}$ and $\{\overline{z}_{L}^{k}\}_{k=1}^{n}$
@@
If the flux distributions and region weights assume their steady-state values such that $\overline{\pi}_{\ell}(dx|k)=\overline{\pi}(dx|k)$ and $\overline{z}_{\ell}^{k}=\overline{z}^{k}$, a global balance condition is satisfied. That is,
@@ IV BAD-NEUS
$0=\sum_{k}\overline{z}^{k}\int\overline{\pi}(dx|k)\mathbbm{E}_{X_{0}=x,J_{0}=k}\left[\sum_{t=0}^{S(1)-1}(f(X_{t},J_{t})-f(X_{t+\tau},J_{t+\tau}))\right].$ (26)
@@
$0=\sum_{k}\sum_{p}c_{\ell}^{p}\overline{z}_{\ell}^{k}\int\phi_{p}(x,k)\overline{\pi}_{\ell}(dx|k)\mathbbm{E}_{X_{0}=x,J_{0}=k}\left[\sum_{t=0}^{S(1)-1}(\phi_{r}(X_{t},J_{t})-\phi_{r}(X_{t+\tau},J_{t+\tau}))\right].$ (28)
@@
$M_{pr}=\sum_{k}\overline{z}_{\ell}^{k}\int\phi_{p}(x,k)\overline{\pi}_{\ell}(dx|k)\mathbbm{E}_{X_{0}=x,J_{0}=k}\left[\sum_{t=0}^{S(1)-1}(\phi_{r}(X_{t},J_{t})-\phi_{r}(X_{t+\tau},J_{t+\tau}))\right].$ (30)
@@
1: Approximate Markov propagator $\tilde{\mathcal{U}}$, Resampling operator $\mathcal{R}_{k}$, starting flux distributions $\{\overline{\pi}_{0}(dx|k)\}_{k=1}^{n}$, initial region weights $\{\overline{z}_{0}^{k}\}_{k=1}^{n}$ with $\sum_{k}\overline{z}_{0}^{k}=1$, basis set $\{\phi_{p}(x,k)\}_{p}$, lag time $\tau$, number of past iterations to retain $h$, total number of iterations $L$, number of strata $n$.
@@
13:return $\{\overline{\pi}_{L}(dx|k)\}_{k=1}^{n}$ and $\{\overline{z}_{L}^{k}\}_{k=1}^{n}$
@@
1: Starting ensemble of walkers $E_{0}=\{(X_{0}^{i},J_{0}^{i},w_{0}^{i})\}_{i=1}^{N},$ with $\sum_{i=1}^{N}w_{0}^{i}=1$, lag time $\tau$, number of past iterations to retain $h$, total number of iterations $L$, number of strata $n$.
@@
1: List of trajectories $T_{\ell}$, lag time $\tau$, basis functions $\{\phi_{p}\}_{p}$, region weights $\{\overline{z}_{\ell}^{k}\}_{k=1}^{n}$
@@
When working with walkers, this mixture distribution corresponds to concatenating the walkers from the last $h$ iterations, and therefore contains more data than using a single iteration. We simply substitute $\overline{\pi}_{\ell,h}$ for $\overline{\pi}_{\ell}$ in all of our algorithms. A unified walker-based algorithm that we use in practice is given in Algorithm 4, with the additional steps needed for NEUS and BAD-NEUS in Algorithm 5.
@@ V.1 Comparison of algorithms
Figure 2: Convergence of the steady-state distribution of the Müller-Brown system for WE as defined in Algorithm 1 (blue, no $\pi$ solve), NEUS as defined in Algorithm 2 (orange, $\pi$ solve using the strata in (37) as a basis), and BAD-NEUS as defined in Algorithm 3 (green, $\pi$ solve using a basis defined by $k$-means clustering, as described in the text).
@@ V.2 Choice of hyperparameters
Having demonstrated that trajectory stratification outperforms WE (without any acceleration strategy) for this system, we now examine the effect of key hyperparameters in NEUS and BAD-NEUS.
@@ V.3 Kinetic statistics
Thus there are $n+m$ total strata, the first $n$ of which contain walkers that last visited $A$, and the remainder of which contain walkers that last visited $B$. The steady-state distribution of walkers which last visited $A$ is proportional to $\pi(x)q_{-}(x)$, and that which last visited $B$ is proportional to $\pi(x)(1-q_{-}(x))$. This ensemble lets us compute dynamical averages. The backward committor projected onto a space of CVs can be computed using
@@ VI Molecular Test System
All molecular dynamics simulations were performed with OpenMM [41]. We compute CVs including backbone RMS deviations (RMSDs) and fractions of native contacts using MDtraj [42]. To enable direct comparison with Ref. 37, we model NTL9(1-39) using the AMBER FF14SB force field [43] and the Hawkins, Cramer and Truhlar generalized Born implicit solvent model [44, 45]. We use a Langevin thermostat with a time step of 2 fs, a temperature of 300. K, and a friction constant of $\gamma=80$ ps-1, which corresponds to the high-viscosity case considered in Ref. 37.
@@
where $Q(X)$ is the fraction of native contacts using the definition of Ref. 46. We determine the native contacts from the crystal structure in the Protein Data Bank (PDB), PDB ID 2HBB [47]. We space 20 $Q_{k}$ values uniformly in the interval $[0.35,0.85]$ and set $\varepsilon_{k}=0.6(Q_{k+1}-Q_{k})$, so that the regions overlap. Thus there are 40 total strata. We define the folded state as full backbone RMSD $<0.28$ nm and $Q>0.85$; the denatured state has full backbone RMSD $>0.80$ nm and $Q<0.35$.
@@
We run each walker until its index process changes and then an additional 10 ps, so the trajectory length and time per iteration are random variables. For the NTL9 simulations, the average trajectory length was 50 ps, and the average time per iteration was 4 $\mu$s. The results in Figure 9 thus required about 204 $\mu$s in total, comparable to the 252 $\mu$s for the haMSM-WE calculation. The time to compute the rate constant by BAD-NEUS and haMSM-WE is much less than the millisecond timescale of a single folding event, which we also expect to be the time to converge WE without any acceleration strategy. BAD-NEUS actually provides some speedup relative to haMSM-WE in the sense that it provides potentials of mean force and committors in its present form, whereas a calculation in the unfolding direction (performed either separately or simultaneously, as here) would be required to obtain these statistics from haMSM-WE. By the same token, we likely could reduce the computation time for BAD-NEUS by “teleporting” walkers from $B$ to $A$ and computing only the forward rate [50]. However, we expect most investigators would seek the mechanistic information provided by the potentials of mean force and committor (which can be used to also obtain reactive currents [24]) if they are investing the computational resources to compute the rate, and free energies computed from ratios of forward and backward rates can serve as checks on the rates [51].
@@
Figure 10: (top) BAD NEUS bin weights for the NTL9(1-39) system. Strata with indices less than 20 correspond to the folding direction (trajectories last in the denatured state), and strata with indices greater than or equal to 20 correspond to the unfolding direction (trajectories last in the native state). (bottom) Deviation of weights ($\ln(\overline{z})$) from those in the last iteration.